A researcher conducted a study of the access speed of 45 hard drives and concluded that his maximum error of estimate was 25. If he were to conduct a second study to reduce the maximum error to 5, about how many hard drives should he include in his new sample?
KCU CHE 214 Laboratory Handout Colligotive Properties 1.) A solution of a sugar in HAcis determined to freeze at 0.25 C below the normal freezing point. What is the molality of this solution? 2.) If the solution contained 0.026 kg of acetic acid, how many moles of the unknown sugar are present? 3.) If the mass of the unknown sugar added to the HAc was 0.303 g. is the sugar a mono- or 4.) Indicate whether the freezing point is located at point A, B, or C in the curve below Cooling Curve for Solution Liquid Solid Sollid Liquid time
Answer #1

Answer to the question)

Given: when number of drives ( N ) = 45

Error ( E) = 25

.

Interpretation of variables used in the formula:

  • value of Z depends of confidence level
  • S = sample standard deviation
  • E = Error
  • N = Sample size

N = (Z*S / E)^2 [formula of sample size]

45 = (Z * S /25)^2

.

on cross-multiplying we get:

(Z *S )^2 = 45 *25^2

(Z*S)^2 = 28125

.

If error is 5 , then size would be as follows:

N' = (Z * S)^2 / 5^2

N'= 28125 /25

N' = 1125 [where N' denotes new sample size]